Changes in Temp:
PART A
The pressure inside a hydrogen-filled container was 2.10 atm at 21 ∘C. What would the pressure be if the container was heated to 91 ∘C ?
Express your answer with the appropriate units.

PART B
At standard temperature and pressure (0 ∘C and 1.00 atm ), 1.00 mol of an ideal gas occupies a volume of 22.4 L. What volume would the same amount of gas occupy at the same pressure and 40 ∘C ?
Express your answer with the appropriate units.
1
Expert's answer
2017-03-17T08:49:35-0400
a) Isochoric process: T1 = 294K, T2 = 384K and P2 = P1*T2/T1 = 2.10atm * 384K/294K = 2.74 atm
b) Isobaric process: T0 = 273K, T1 = 313K and V1 = V0*T1/T0 = 22.4L * 313K/273K = 25.7 L